OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $8.55. If the number of dimes was tripled and the number of nickels was decreased by 2, the value of the coins would be $23.45. How many nickels and dimes does he have?

OpenStudy (amistre64):

.05n + .10d = 8.55 .05(n-2) + .10(3d) = 23.45

OpenStudy (amistre64):

5n + 10d = 855 5(n-2) + 10(3d) = 2345 5n + 10d = 855 5n -10 + 30d = 2345 5n + 10d = 855 5n + 30d = 2355 d = (2345(5)-855(5))/(150-50) = 74.5 hmmm

OpenStudy (amistre64):

oh , i typoed in a 2355 lol

OpenStudy (amistre64):

75 dimes

OpenStudy (amistre64):

8.55 -7.50 ----- 1.05 = 21 nickels right?